Home‎ > ‎Reports‎ > ‎crew113‎ > ‎

day13


Crew #113 Log Book for February 23, 2012

Photos of the day





Commander's Check-In Report

Date and time: February 23 2012, 19:15

Written by: Alexandre Mangeot, Crew 113, Commander

Crew Physical Status: Ok

Time Departed/Returned from EVA: 5PM -> 6:15PM

Brief Narrative of Field Mission Results: Sample from site 1 to 4 collected. The samples will be analyzed by crew 114 and the remote science team.

EVA Data/Interpretations: N/A

Engineering/Hab Maintenance: External water tank is empty, resupply arranged with DG.

Packages delivered to the Hab: N/A

Report Transmission Schedule: 8:30PM

Plans for Tomorrow: Sample soil of sites 8 to 10

Inventory:

Support Requested:

Miscellaneous:

Commander's Report

Date and time: February 23 2012, 19:00PM

Written by: Alexandre Mangeot, Crew 113, Commander

One of the past crew members asked me during the night if the crewmember are willing to make a summary of their work for compilation and preparation of a scientific paper for a incoming symposium.

Briefing at 8:00. The wind was too strong to allow an EVA. Only Alexandre and Baptiste have been outside for the engineer rounds.
I also asked to my crew to prepare a summary of their work here in order to comply with the demands for the scientific paper.

Dr Robert Zubrin send me an email this morning asking me to send the final report of the mission at Friday night. It is an opportunity to begin a summary of the work for The Mars Society final report and for the scientific paper as well.

I begun received the reports from my crewmates at 10:00. Some of them had to be rewritten because of the formulation (1st or 3rd person narrative text instead of passive form). I finished compilation of the summaries at noon. I have sent the final report version beta for reviewing, correction and comment by my crewmembers and our remote director as well.

Some lines must be modified to take into account the activities of today and tomorrow. But since the major part of the report is already done, we should be able to send the final report on time.

After lunch, we rested for 1 and a half hours. When I woke up, the wind has calmed down and I decided to go for EVA with Valentina and Pieter. We performed the sampling requested by the Crew 114 remote science team at site 1 to 4. We came back at sunset.

I went Out of Sim for a couple of pictures of the Hab lighted by the last sun rays.

EVA Report

Date and time: February 23 2012, 19:16

Written by: Pieterjan Grandry, Crew-113, Support

Time: 17:00 - 18:15

Crew Members: Pieterjan Grandry, Alexander Mangeot, Valentina Karga

Site Location: Near the Hab

Transit Mode: ATV, foot

Objectives: Take samples according to the protocol.

Results: We located the three spots and managed to take samples from the soil according to the protocols, only the lighter couldn't burn the tools in order to sterilize them, so we just cleaned them.

Lessons Learned: Bring good lighter when you need to use one in the wind.

Engineering Report

Date and time: 23 February, 2012, 18:30
Written by Alexandre RICHARD, Crew 113, Crew Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): <0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5

Oil:
ATV Oil Quantity (# of Quarts in storage): 20
Car Oil Quantity (# of Quarts in storage): 40

Water Status:
Outside Potable Water Tank Level (inches from bottom): 12
Trailer Potable Water Tank Level (inches from bottom): 15
Hab Potable Water Tank Level (inches from bottom): EMPTY
Potable Water Meter Reading: 0051307.5
Into GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): NO
Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): YES: About 33%
Tire Status: Good
Odometer at end of day: 126234.2
Notes/Comments on Hab Car:N/A

ATV 1 (Opportunity):
Used: YES
Oil Checked: YES: Full
Fuel Consumed: 1 gal
Tire Status: GOOD
Today's Run Time: 00:45
Season's Run Time: 86:55
Comments:-

ATV 2 (Spirit):
Used: YES
Oil Checked: YES: Full
Fuel Consumed: 1 gal
Tire Status: GOOD
Today's Run Time: 00:45
Season's Run Time: 81:55
Comments:-

ATV 3 (Viking I):
Used: YES
Oil Checked: YES: Full
Fuel Consumed: 1 gal
Tire Status: GOOD
Today's Run Time: 00:45
Season's Run Time: 84:55
Comments:-

Heating and Ventilation:
Thermostat setting upstairs (Farenheit):65°F
Actual temperature upstairs (Farenheit):64°F

Thermostat setting downstairs (Farenheit):65°F
Actual temperature downstairs (Farenheit):61°F

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:
N/A

General Engineering Notes/Comments: N/A

Chef's Report

Date: February 23 2012
Written by: Moeglin Baptiste

1a) Today was

- a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) or activities were prepared to mark the celebration?

*No*

2) List the foods served at today's main meal (usually the evening meal),
giving full names of each.

creamy potato soup + spaghetti + beef and tomatoes

3) List any main meal foods not finished by the crew (leftovers stored for
later use or discarded)

spaghetti (discarded)

4) Recipes for anything more complicated than following package
directions exactly, or rehydrating.

5) Lessons learned (if any)

6) Comments/questions for Kim and Jean, the food study investigators
*No*